OpenStudy (anonymous):

Hello, i am terribly stuck on this question could someone please help me.I know these two point on a line, they are (2a,2c) and (2b,0) it asks me then to find the equation of the perpendicular bisector of this line. Ive spent 40 minutes on it and i havent gotten anywhere. Ive got up to the slop of the perpendicuar bisector and 1 point (the midpoint of AB (the line between coordinates)) Thanks to any help i can get

OpenStudy (anonymous):

find the mindpoint. then find the slope through the two points. slope = m slope of the perpendicular is -1/m, where m is the slope through the two points. use point slope form to get the equation of the perpendicular bisector.

OpenStudy (dumbcow):

First get slope of line: slope = 2c/(2a-2b) = c/(a-b) Next get midpoint because "bisector" means the line will have to intersect at midpoint midpoint = ((2a+2b)/2 , 2c/2 ) = (a+b , c) slope of perpendicular line is opposite reciprocal of original slope --> -(a+b)/c next get the equation of line using above slope and the midpoint y-c = -(a+b)/c (x -(a+b)) y = -(a+b)/c x + (a^2-b^2)/c + c y = -(a+b)/c x + (a^2 -b^2 +c^2)/c

OpenStudy (dumbcow):

im sorry theres a typo, perpendicular slope should be -(a-b)/c

OpenStudy (anonymous):

THANK YOU SOOOOOOOOOO MUCH!! saved my evening

OpenStudy (dumbcow):

haha your welcome